OpenStudy (anonymous):

Assume the $10,000 Treasury bill, 4% for 13 weeks. Calculate the effective rate of interest.(Use calendar year. Round your answer to 2 decimal places. Omit the "%" sign in your response.)

OpenStudy (anonymous):

Sorry but no I can't. Effective Rate is actually where I am having the Most trouble??

OpenStudy (anonymous):

I can see above where you have figured out the Proceeds, but I do not see how to figure out the Effective Rate??

OpenStudy (anonymous):

I am sorry, but I just do not follow how to do this??? What exactly do you need me to post?

OpenStudy (anonymous):

Do you understand the below formula? Effective rate of interest = interest / proceed * time

OpenStudy (anonymous):

Not really? Sorry!

OpenStudy (anonymous):

Do you know what you're looking for in your post?

OpenStudy (anonymous):

Assume the $10,000 Treasury bill, 4% for 13 weeks. Calculate the effective rate of interest.(Use calendar year. Round your answer to 2 decimal places. Omit the "%" sign in your response.) Effective rate of interest %

OpenStudy (anonymous):

This is the exact way it is posted on my homework site. I copied and pasted it.

OpenStudy (anonymous):

No, don't just post the question! Do you know what the question asking?

OpenStudy (anonymous):

Its asking for the Effective rate of intrest as a %

OpenStudy (anonymous):

YES, Effective rate of intrest !

OpenStudy (anonymous):

Have you learned the formula to find Effective rate of interest?

OpenStudy (anonymous):

No, this is an online class and I have not been in any acounting classes before. I am also 43 years old and have never had an online class til now, and I am only in my 2nd semester of colleuge ever.

OpenStudy (anonymous):

This is the formula I've provided: Effective rate of interest = interest / ( Proceed * Time )

OpenStudy (anonymous):

Ok, I will try to figure it out. I apolgize for taking up so much of your time. Thank You once again!!!

OpenStudy (anonymous):

Did you see the interest I've worked out?

OpenStudy (anonymous):

All you do is plug the numbers into your calculator!

OpenStudy (anonymous):

Is that the $100.00 at the top of our posts??

OpenStudy (anonymous):

Interest = p * r * t = 10,000 * .04 * 13/52 = $100

OpenStudy (anonymous):

Yes, that part I understand.

OpenStudy (anonymous):

Why do you ask the interest is 10,000 ???

OpenStudy (anonymous):

Do you see the Proceed I also have worked it out ?

OpenStudy (anonymous):

The Proceed is the $9900. correct?

OpenStudy (anonymous):

All you do is to read: Proceed = 10,000 - 100 = $9,900

OpenStudy (anonymous):

Can you plug these worked out result into: Effective rate of interest = interest / ( proceed * time )

OpenStudy (anonymous):

I am sorry but I have to go, someone has just stopped over, and I am also feeling a bit stupid at this moment, I will go to the school tudor and perhaps I will get it then. Thanks again for all of your efforts and help!!!!!!

OpenStudy (anonymous):

Okie, all you do is to plug the number in = 4.04